What is the probability for a leap year to have 52 Monday’s and 53 Sunday’s.

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
:  In a leap year, there are 366 days. Out of these 364 days make 52 weeks and hence 52 saturday. Remaining 2 days can be: Sun-Mon, Tue-Wed, Wed-Thu, Fri-Sat, Sat-Sun.
Total number of possible outcomes = 7
No of saturday outcomes = 2 (Fri-Sat, Sat-Sun)
Then probability of 53 sundays = 2/7
 
As no period of time has been specified, it can be taken that the extra two days of a leap year can be any two days of the week. The probability that one of them is a Monday is 2/7. The probability that one of them is not a Monday is 1 - 2/7 = 5/7.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)